### v1

# Noise-augmented Chaotic Ising Machines for Combinatorial Optimization and Sampling

Abstract
The rise of domain-specific computing has led to great interest in Ising machines, dedicated hardware accelerators tailored to solve combinatorial optimization and probabilistic sampling problems. A key element of Ising machines is stochasticity, which enables a wide exploration of configurations, thereby helping avoid local minima. Here, we evaluate and improve the previously proposed concept of coupled chaotic bits (c-bits) that operate without any explicit stochasticity. We show that augmenting chaotic bits with stochasticity leads to better algorithmic scaling in combinatorial optimization problems, comparable to the performance of probabilistic bits (p-bits) which have explicit randomness in their update rules. We first demonstrate that c-bits surprisingly follow the quantum Boltzmann law in a 1D transverse field Ising model, despite the lack of explicit randomness. We then show that c-bits exhibit critical dynamics similar to those of stochastic p-bits in 2D Ising and 3D spin glass models, with promising potential to solve challenging optimization problems. Finally, we propose a noise-augmented version of coupled c-bits via the powerful adaptive parallel tempering algorithm (APT). The noise-augmented c-bit algorithm outperforms fully deterministic c-bits running versions of the simulated annealing algorithm. Chaotic Ising machines closely resemble coupled oscillator-based Ising machines, as both schemes exploit nonlinear dynamics for computation. Oscillator-based Ising machines may greatly benefit from our proposed algorithm, which runs replicas at constant temperature, eliminating the need to globally modulate coupling strengths. Mixing stochasticity with deterministic c-bits creates a powerful hybrid computing scheme that can bring benefits in scaled, asynchronous, and massively parallel hardware implementations.
I.   Introduction
s Refer to caption
Fig. 1: (a) A probabilistic bit (p-bit) contains a random number generator with uniform distribution between -1 and 1. $\tanh(I_{i})$ is a threshold for latching $m_{i}$. (b) A chaotic bit (c-bit) features deterministic billiard dynamics with a tunable slope. The billiard is periodic, given that $I_{i}$ is held constant. When the billiard reaches $-1$ or $+1$, $m_{i}$ is latched, and the billiard changes direction. (c) p-bits and c-bits form similar, asynchronous architectures, employing the same synaptic function. (d) A 20-bit factorizer conceptualized with p-bits, represented as a graph with 310 nodes and 1200 edges. (e) The 3D spin glass problem, with coupling strengths $J_{ij}\in\{-1,+1\}$. (f) The adaptive parallel tempering (APT) algorithm. Replicas of the same network run at different inverse temperature $\beta$ in parallel. Swaps attempts are made between pairs of replicas with the closest $\beta$, according to the Metropolis criterion (Eq. 15).
As Moore’s law stagnates, there is growing interest in unconventional computing schemes and hardware accelerators. Among these emerging technologies are Ising Machines [1], employing devices such as probabilistic bits (p-bits) [2, 3, 4, 5], coupled oscillators [6, 7, 8, 9, 10, 11, 12], photonic devices [13, 14, 15, 16], and other nonlinear elements [17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Ising machines have traditionally focused on providing scaling and prefactor improvements in solving combinatorial optimization problems [27, 28, 29]. Efforts to link Ising machines to generative machine learning [30, 31, 32, 33, 34] and quantum simulation problems are also underway [35, 36, 37]. Here, we evaluate and improve an earlier Ising machine concept based on chaotic bits (c-bits), originally proposed in Ref. [38]. Chaotic bits resemble coupled oscillators, following a set of deterministic neuron update rules without explicit randomness. However, unlike oscillator-based Ising Machines that encode the Ising spin in the continuous phase of oscillators, c-bits make use of oscillating billiard balls that set a latch to a $+1$ or $-1$ state. As such, chaotic bits do not require cumbersome subharmonic injection locking schemes to binarize naturally continuous phase variables.
The potential hardware implementation of chaotic Ising machines without any explicit random number generators, while retaining similarity to the mathematics of p-bits, is appealing. However, as we demonstrate in this paper, c-bits without any randomness cannot straightforwardly employ the most powerful Monte Carlo algorithms, such as Adaptive Parallel Tempering, which offer better algorithmic scaling for the Circuit SAT and 3D Spin Glass problems. We propose and evaluate a hybrid computing scheme wherein deterministic c-bits are augmented with stochasticity for improved performance in combinatorial optimization and sampling. The approach we propose may be applied to similar devices, such as coupled oscillators, where a continuous global modulation of analog coupling elements may be inconvenient.
Probabilistic computers perform a discrete Markov Chain Monte Carlo (MCMC) algorithm, simulated in software via Gibbs sampling (FIG. 1c). p-bit networks are typically described by second-order interactions in energy and stochastic neurons:
$m_{i}=\text{sgn}[\tanh(\beta I_{i})-\text{rand}_{u}(-1,1)]$ (1)
$I_{i}=\sum_{j}J_{ij}m_{j}+h_{i}$ (2)
where $m_{i}\in\{-1,+1\}$ and $\text{rand}_{u}(-1,1)$ is a random uniform distribution between $-1$ and $1$ (FIG. 1a, c). $\beta$ is inverse temperature. $J_{ij}$ and $h_{i}$ represent weights and biases. At equilibrium, p-bit networks sample from the Boltzmann distribution, given by:
$p(\{m\})=\frac{1}{Z}\exp[-\beta E(\{m\})]$ (3)
$E(\{m\})=-\frac{1}{2}\sum_{i,j}J_{ij}m_{i}m_{j}-\sum_{i}h_{i}m_{i}$ (4)
where $\{m\}$ is a spin configuration, $Z$ is the partition function, and $E(\{m\})$ is the energy of a spin configuration. Chaotic bits can be formulated to follow similar equations as p-bits, that describe an oscillating billiard ball and a latch:
$\frac{dx_{i}}{dt}$ $=-m_{i}+\tanh(\beta I_{i})$ (5)
$x_{i}=+1$ $\implies\text{set }m_{i}\text{ to }+1$
$x_{i}=-1$ $\implies\text{set }m_{i}\text{ to }-1$
where $x_{i}$ represents the position of a billiard between the boundaries $x_{i}=-1$ and $x_{i}=+1$. The billiard has a tunable slope, ${dx_{i}}/{dt}$, which is dependent on the input $I_{i}$. When the billiard $x_{i}$ reaches $-1$ or $+1$, the c-bit state $m_{i}$ is latched to that value, and the billiard changes direction (FIG. 1b). Note an interesting parallel to p-bits. In Gibbs sampling, simultaneously updating connected nodes leads to pathological oscillations [39, 40]. For c-bits, the differential equation formulation in continuous time prevents simultaneous updates naturally, as long as the time difference between subsequent billiard ball arrivals is greater than the time it takes to update a latch. When c-bits are discretized in software, latches are set instantaneously within a single time step, $\Delta t$, chosen to solve the differential equations. In Section VII, we briefly discuss implications of this update mechanism for eventual hardware implementations. It is important to note that p-bits and c-bits have different notions of time. There is no clear correspondence between a p-bit Monte Carlo sweep (MCS), in which Eqs. 1 and 2 are sequentially solved for each p-bit in the network, and the time variable $t$ of a c-bit network. In this study, we sample the state $\{m\}$ of a continuously updating c-bit network at natural numbers $t=\{0,1,2,\ldots,t_{a}\}$.
Individual p-bits in a system observe the following relationship:
$\frac{p(m_{i}=+1)}{p(m_{i}=-1)}=\exp{(2\beta I_{i})}$ (6)
Intuitively, c-bits try to achieve the same by modifying the lifetime of up and down trajectories of the billiard balls, conditioned on the state of their neighbors (FIG. 1b):
$\frac{\tau(m_{i}=+1)}{\tau(m_{i}=-1)}=\frac{\left|\displaystyle\frac{dx_{i}}{dt}\quad\mbox{when $m_{i}=-1$}\right|}{\left|\displaystyle\frac{dx_{i}}{dt}\quad\mbox{when $m_{i}=+1$}\right|}=\exp{(2\beta I_{i})}$ (7)
where $\tau$ is the lifetime of a c-bit at its current state. The key point is that $\tau$ are deterministic, unlike a p-bit whose state is stochastically sampled at each update. Eq. 5 is not the only valid c-bit definition [41]. As long as the ratio in Eq. 6 is observed, a c-bit could be defined using different velocities.
This intuitive explanation does not show that c-bit networks sample from the Boltzmann distribution as stochastic p-bits do. Theoretical work has shown convergence to the Boltzmann distribution for simple toy models [42]. Experimental work suggests that c-bit networks seem to closely approximate the Boltzmann distribution for 2D Ising models of small sizes [41]. Furthermore, it has been shown that c-bits match stochastic p-bits for equilibrium statistics on the 2D ferromagnetic Ising model [38] and the Potts model [41]. c-bits have been used to determine the critical temperature ($T_{c}$) and critical exponents ($\nu$, $\beta$, and $\gamma$) of the Ising universality class [41]. Simulated annealing has been demonstrated with c-bits [38, 43]. Additionally, CMOS and analog VLSI implementations have been proposed [44, 45]. Nonetheless, the theoretical equivalence of c-bits to p-bits without explicit noise is far from clear. Moreover, the initial phase randomization and rounding errors in the solution of ODEs (or physical noise in real systems) might be responsible for this striking correspondence.
Here, we subject c-bits to more stringent tests on three important problem classes, comparing their performance to p-bits. These tests involve a quantum problem, criticality in 3D spin-glasses, and hard combinatorial optimization in the form of integer factorization with invertible logic gates. We conclude that while c-bits may perform similarly to p-bits in algorithms such as simulated annealing, they may benefit from added stochasticity, either through their update rules (for example, by adding jitter to the billiard periods), or by adding explicit random swaps used in powerful tempering algorithms.
Most existing work employs a definition of ${dx_{i}}/{dt}$ that grows exponentially at cold temperatures ($\beta\gg 1$), with the exception of one previous study [41]. This is problematic because optimization algorithms such as simulated annealing and adaptive parallel tempering require cold temperatures which may result in unbounded slopes. In this study, we employ a c-bit definition that has a maximum billiard speed of 2 ($\sum\tau(m_{i}=\pm 1)=2$, see, Eq. 5), suitable for software and hardware implementation. Unless otherwise specified, we employ the Euler’s Method with a step of $\Delta t=0.1$ to simulate the dynamics of coupled c-bits.
Our bounded c-bit choice is also a potential improvement over previous hardware implementations, which may be suffering from accuracy losses due to exponentially growing slopes. For example, the anomalous increase of energy plots in a prior FPGA implementation is likely due to the overflows and underflows in slopes ([43]).
Finally, in all of our experiments, the initial phases of the billiard balls and the states of the latches are randomized. The asynchronous dynamics of the billiard balls, coupled with this phase randomization ensures nonsimultaneous updates with quasi-random arrivals. In addition, rounding errors and the choice of a naive integration scheme provide additional noise in our simulations. We believe these seem to be the key reasons behind the near-equivalence of c-bits to p-bits where explicit and implicit inclusion of noise seems to play a key role.
II.   Sampling from the 1D Transverse Field Ising Hamiltonian with c-bits
Fig. 2: Chaotic bits emulate a 1D ferromagnetic chain ($J_{ij}=+2$) of 8 qubits described by the quantum Transverse Ising Hamiltonian (Eq. 8). For emulation purposes, we use Suzuki-Trotter decomposition with $R=250$ replicas. We show average magnetization as a function of the transverse field ($\Gamma_{x}$) when the system is at constant inverse temperature $\beta=10$. A symmetry breaking magnetic field of $h_{i}=1$ is applied in the $+\hat{z}$ direction such that when $\Gamma_{x}=0$, all spins point in the $+\hat{z}$ direction. Red dots show the results of 80 independent trials of c-bit simulations, and a single set of spin configurations is recorded at the end of $t_{a}=1000$ time steps, for each trial. Blue triangles show the average numerical result. The green dashed line is an analytical solution from solving Eq. 11 as a function of $\Gamma_{x}$.
As a difficult sampling problem that requires high-quality pseudorandom number generators [46], we consider a 1D ferromagnetic chain of 8 qubits at constant inverse temperature $\beta=10$. This system is described by the transverse Ising Hamiltonian in 1D,
$H_{\text{Q}}=-\sum_{i<j}{J_{ij}\sigma^{z}_{i}\sigma^{z}_{j}}-\sum_{i}{h_{i}\sigma^{z}_{i}}-\Gamma_{x}\sum_{i}{\sigma^{x}_{i}}$ (8)
where $\sigma_{i}^{x}$ and $\sigma_{i}^{z}$ are Pauli spin matrices at site $i$. The interaction weight $J_{ij}=+2$ for nearest neighbors (with a periodic boundary condition) and $J_{ij}=0$ otherwise. $h_{i}=+1$, representing a symmetry breaking magnetic field in the $+\hat{z}$ direction. The Suzuki-Trotter transformation, [47, 35] maps the 1D quantum Hamiltonian to the following 2D classical Hamiltonian:
$H_{\text{C}}=-\sum\limits_{k=1}^{R}\left[\sum\limits_{i<j}\cfrac{{J}_{ij}}{R}{m}_{i,k}{m}_{j,k}+\sum\limits_{i}{\cfrac{h_{i}}{R}\,m_{i,k}}\right.$
$\left.+\sum\limits_{i}{J}_{\perp}{m}_{i,k}{m}_{i,k+1}\right]$ (9)
where $J_{\perp}=-{1}/{2\beta}\ln{\tanh{\left({\beta\Gamma_{x}}/{R}\right)}}$. Note the unusual dependence of this coupling to the inverse temperature, $\beta$.
For an infinite number of replicas ($R\rightarrow\infty$), this quantum-to-classical mapping is exact. In practice, the error scales as $\mathcal{O}(1/R^{2})$ for a given $\beta$. Here, we employ $250$ replicas, thus our system contains $N=8$ qubits and $R=250$ replicas, totaling $2000$ classical spins. We simulate our classical system using c-bits. As we vary the value of the transverse field $\Gamma_{x}$, we measure the average magnetization:
$\langle m_{z}\rangle=\cfrac{1}{NR}\sum_{k=1}^{R}{\sum_{i=1}^{N}{m_{i,k}}}$ (10)
In theory, the system should obey the quantum Boltzmann Law:
$\langle m_{z}\rangle=\cfrac{\text{tr}\left[\exp{(-\beta H_{\text{Q}})}\sum_{i}{\sigma_{i}^{z}}\right]}{\text{tr}\left[\exp{(-\beta H_{\text{Q}})}\right]}$ (11)
FIG. 2 shows that numerical results from c-bits show strong agreement with the quantum Boltzmann law, obtained by solving Eq. 11 as a function of $\Gamma_{x}$. A similar result has been shown in a previous work using stochastic p-bits [48]. We observe that c-bits appear to show comparable performance to p-bits, despite the fact that c-bits only use explicit randomization for the initialization of spins and phases at the start of each simulation.
III.   Critical Scaling Dynamics: c-bits vs p-bits
Fig. 3: (a) Residual energy density at the critical point ($\rho_{E}^{c}$) as a function of annealing time for a $50\times 50$ ferromagnetic Ising model lattice with nonperiodic boundary conditions ($J_{ij}=+1$). We perform linear regression on the 5 leftmost points of each data series to avoid finite-size effects. Error bars show $95\%$ bootstrap confidence intervals. (b) $\rho_{E}^{c}$ vs. annealing time for cubic spin-glass instances ($L=11$, $J_{ij}\in\{-1,+1\}$). (c) $\rho_{E}^{c}$ as a function of inverse annealing velocity for the 2D ferromagnetic Ising model, using p-bits. Different annealing schedules are shown in a data collapse by considering the quench rate at the critical point, $\alpha$. (d) $\rho_{E}^{c}$ vs. inverse annealing velocity for the 2D ferromagnetic Ising model, using c-bits.
Next, we discuss critical dynamics of c-bit networks for the 2D Ising model and the 3D spin glass problem. The idea is to quench c-bit networks at varying anneal rates and measure the residual energy as a proxy of topological defects. The density of these defects are qualitatively predicted by the Kibble-Zurek Mechanism (KZM) [49], which relates equilibrium correlation lengths and relaxation times to defect densities borne out of non-equilibrium dynamics. In the present context, the study of KZM allows a comparison of c-bit performance with that of p-bits. In addition, residual energy may be a useful figure of merit for combinatorial optimization where scaling comparisons between different algorithms can be made [50]. The residual energy density at the critical point is defined as:
$\rho_{E}^{c}=\frac{\langle E-E_{c}\rangle}{n}$ (12)
where $E_{c}$ is equilibrium energy at the critical temperature and $n$ is the number of spins in the system. From the theory of phase transitions, phenomenological scaling arguments can be made to describe the residual energy in terms of an anneal with varying velocities (see Appendix .1 for details):
$\rho_{E}^{c}\propto\alpha^{\kappa_{c}}$ (13)
where $\kappa_{c}$ is the exponent describing the scaling behavior of the residual energy density at the critical point, and $\alpha$ is the velocity of the anneal measured at the critical temperature for any schedule that starts from some initial temperature and ends at the critical temperature:
$\alpha=\left.\frac{d\lambda(t)}{dt}\right|_{t=t_{a}}$ (14)
One way to make contact with known critical exponents and our numerically calculated exponents is to measure the number of defects via kinks [51] in Ising models, but our purpose here is to provide a comparative analysis with p-bits and c-bits, so direct verification of the exponents is not necessary.
First, we consider a $50\times 50$ ferromagnetic Ising model lattice with nonperiodic boundary conditions ($J_{ij}=+1$). For a given annealing time $t_{a}$, $\beta$ is geometrically increased from $\beta=0.001$ to the critical temperature, $\beta={T_{c}}^{-1}$. For this problem, we use the exact critical temperature calculated at the thermodynamic limit [52] as an approximate $T_{c}$ for the finite-size model: $T_{c}={2}/{\ln(1+\sqrt{2})}\approx 2.269$). When annealing terminates at the critical point, the energy of the spin configuration is recorded to calculate $\rho_{E}^{c}$. We average over 3000 randomized trials.
Although there is no obvious correspondence between a Monte Carlo sweep for behavioral models of p-bits and the time step of a c-bit, power-law exponents can be compared objectively since they do not depend on prefactors. FIG. 3a shows that c-bits and p-bits demonstrate similar critical scaling dynamics on the 2D ferromagnetic Ising model. Linear regression is performed on the 5 leftmost points in order to avoid finite-size effects. Both p-bit and c-bit curves exhibit downward concavity, but different time scales mean that this concavity does not manifest at the same time $t_{a}$ for p-bits and c-bits. We provide the results of using Euler’s method in steps of $\Delta t=0.1$ and $\Delta t=0.01$. The c-bit curve using $\Delta t=0.01$ appears to follow a similar power law scaling, but with a steeper slope $|\kappa_{c}|$. The reason for this discrepancy is not clear. On the one hand, smaller $\Delta t$ implies higher theoretical precision, while on the other hand, smaller $\Delta t$ leads to more round-off errors due to double precision. It must also be noted that a lower $\Delta t$ requires more computational effort, and in hardware, $\Delta t$ may have a lower bound dictated by the physical latch required to set the state of the oscillator.
FIG. 3b shows a similar experiment conducted on select instances of the 3D spin glass problem, studied by D-Wave [50]. We consider 300 different spin glass instances on a cubic lattice with side length $L=11$ and $J_{ij}\in\{-1,+1\}$. Analogous to FIG. 3a, for a given annealing time $t_{a}$, we geometrically increase $\beta$ to the inverse critical temperature $1/T_{c}$. For 3D spin glasses, $T_{c}$ is approximately calculated as $\approx 1.1$ [53]). For each of 300 problem instances, we take an ensemble average over 10 random seeds. We again observe similar power-law scaling. Our p-bit results match the numerical results in the literature, with $\kappa_{c}\approx 0.51$ [50]. However, the values of $\kappa_{c}$ do not seem to be similar when comparing between p-bits and c-bits. The $\Delta t$ dependence of c-bit slopes also is an indication that the underlying dynamics for c-bits might depend on separating near arrivals that latch a state (where Eq. 2 is always assumed to be infinitely fast). We conclude that while c-bits seem to obey a similar power law as p-bits, their dynamics seems different in subtle ways.
In FIG. 3c-d, we show how both c-bits and p-bits exhibit a clear power-law relationship with critical exponents by comparing different annealing schedules with different initial temperatures. As Eq. 13 predicts, the simulations should exhibit universal behavior if plotted as a function of how fast the anneal is performed near the critical point. When plotted as function of annealing velocity, $\alpha$ (Eq. 14), all c-bit and p-bit plots fall on top of each other, demonstrating an excellent collapse. The different exponents of c-bits and p-bits and the time-step dependence of the c-bit networks indicate that c-bit networks may not be a drop-in replacement for p-bit networks whose steady-state provably takes samples from the Boltzmann distribution [54].
IV.   Adaptive Parallel Tempering
Input: weights $J_{ij}$, biases $h_{i}$, number of replicas $N$, inverse temperature profile $\beta$, time steps per swap, simulation time $t_{a}$
Output: Spin states $\{m\}$ corresponding to the minimum energy $E_{0}$
Initialize $N$ system replicas to random spins;
for time $t=0$ to $t_{a}$ do
       for $N$ replicas at different $\beta$ do
             Simulate p-bit Monte Carlo sweeps using Eq. 1 and 2, or c-bit dynamics using Eq. 2 and 5;
       end for
      if time $t$ is a multiple of time steps per swap then
             // Probabilistic swap attempt
             if it is an even numbered swap attempt then
                   Select (even, odd) replica pairs;
            else
                   Select (odd, even) replica pairs;
             end if
            for each replica do
                   Compute energy $E\{m\}$ (Eq. 4);
             end for
            for each pair of adjacent replicas do
                   if the Metropolis criterion is met (Eq. 15) then
                         Swap the spin state $\{m\}$ between replicas;
                         // For c-bits, also swap the billiard state $\{x\}$ between replicas
                   end if
             end for
       end if
end for
Algorithm 1 Adaptive Parallel Tempering with p-bits or c-bits
We turn our attention to optimization. Critical scaling dynamics are intriguing from a physics perspective, but it is not essential that c-bits obey the same physics as p-bits in an optimization context. For optimization, we use an adaptive version of the parallel tempering algorithm, abbreviated as APT [55, 56, 57]. A given p-bit network is duplicated into $N$ system replicas, with each replica at a different inverse temperature $\beta$ (FIG. 1f). In fixed time intervals (time steps per swap attempt), a probabilistic swap attempt is conducted between adjacent replica pairs according to the Metropolis criterion:
$P(\text{swap})=\min[1,\exp(-\Delta\beta\Delta E)]$ (15)
If this condition is met, then the adjacent replicas swap their spin states $\{m\}$. Alternatively, replicas may swap their temperature values with a rearrangement of replica indices, which may be more convenient in dedicated hardware implementations. Intuitively, hot replicas (small $\beta$) explore the state space of possible spin states $\{m\}$. The Metropolis criterion makes low energy configurations more likely to swap to cold replicas (high $\beta$). Over time, the coldest system replica tends toward the ground energy, $E_{0}$.
The APT algorithm can be applied using either p-bits or c-bits. While the neuron update rule is stochastic for p-bits (Eq. 1), it is deterministic for c-bits (Eq. 5). Replica swaps via the Metropolis criterion are always probabilistic (Eq. 15). Therefore, while p-bit APT is a purely probabilistic algorithm, c-bit APT is a hybrid chaotic-probabilistic scheme, where neuron updates are deterministic and replica swap attempts are probabilistic. In typical p-bit APT, the great majority of the random numbers are used for neuron updates rather than replica swap attempts. As such, if c-bit APT proves to be viable, then it would require orders of magnitude fewer random numbers. However, this potential benefit critically depends on the hardware cost of implementing c-bits, which may be greater than that of p-bits. Given the relative ease of creating pseudo random number generators through compact linear-feedback-shift-register circuits [58], we suspect that the c-bits may surreptitiously contain PRNG-like circuits in their implementation. Our view is that the p-bits and c-bits are very similar in their computational power, but p-bits enjoy a mathematical certainty that they sample from the Boltzmann distribution at steady state. We further discuss these ideas in Section VII.
The algorithm is referred to as ‘adaptive’ because we employ an instance-specific pre-processing method that finds a suitable $\beta$ schedule for a given problem. As shown in FIG. 5d, this process ensures that replica-swap probabilities are nearly uniform, avoiding bottlenecks during the exchange process (see Appendix .2). The complete APT algorithm is summarized in pseudo code (Algorithm 1).
V.   Optimization of the 3D Spin Glass Problem
Fig. 4: Residual energy density ($\rho_{E}^{f}$) as a function of total computational time ($Nt_{a}$) for cubic spin-glass instances with side length $L=7$ ($J_{ij}\in\{-1,+1\}$). Curves scale as a power law: $\rho_{E}^{f}\propto t_{a}^{-\kappa_{f}}$. Error bars show 95% bootstrap confidence intervals.
We consider 300 spin glass instances on a cubic lattice with side length $L=7$ and $J_{ij}\in\{-1,+1\}$, borrowed from Ref. [50]. Purely probabilistic APT (p-bits) and hybrid chaotic-probabilistic APT (c-bits) are performed using similar parameters. We conduct APT pre-processing on 1 spin glass instance, returning a $\beta$ schedule for 14 system replicas. In principle, to achieve the best performance, one could apply the APT pre-processing to each of 300 different problem instances and obtain 300 unique $\beta$ schedules. Since this is a comparative study between p-bits and c-bits, we forego this step. Tuning parameters is not of particular concern.
For a given simulation time $t_{a}$, each of 300 spin glass instances is optimized using APT. For each simulation, we record the lowest energy $E$ sampled across all replicas. We introduce residual energy as an optimization performance metric, defined as:
$\rho_{E}^{f}=\frac{\langle E-E_{0}\rangle}{n}$ (16)
where $E_{0}$ is the ground state energy and $n$ is the number of spins in the system. FIG. 4 shows $\rho_{E}^{f}$ vs. computational time, where the time axis accounts for the total time steps summed across all $N=14$ replicas that run in parallel. We take an ensemble average over 40 random seeds that randomize initial states (for p-bits and c-bits) and phases (for c-bits).
We devise a similar experiment for simulated annealing. In order to draw a comparison to APT, simulated annealing is conducted 14 times in parallel, with $\beta$ increasing linearly from 0 to 10 during simulation time $t_{a}$. We record the lowest energy sampled across all replicas to compute the residual energy $\rho_{E}^{f}$. We take an ensemble average over 40 random seeds.
To compare p-bit and c-bit performance, we employ similar power-law scaling arguments as we did for critical scaling dynamics. $\rho_{E}^{f}=At_{a}^{-\kappa_{f}}$, where $A$ is a constant accounting for the difference in p-bit and c-bit prefactors, and $\kappa_{f}$ is the power law scaling exponent. FIG. 4 shows that p-bit SA exhibits a similar scaling exponent $\kappa_{f}$ to c-bit SA, while purely probabilistic APT yields a similar $\kappa_{f}$ as hybrid probabilistic-chaotic APT. Furthermore, the APT algorithms show $\kappa_{f}$ of larger magnitude than the SA algorithms, indicating a faster convergence to the ground energy $E_{0}$. Two conclusions can be drawn from our results. First, in the optimization setting, there is a a remarkable similarity between c-bits and p-bits since they essentially show the same scaling for both SA and APT. Second, the hybrid APT algorithm we propose for c-bits shows superior performance over its fully deterministic counterpart. Besides better algorithmic scaling, the APT algorithm we propose enjoys another benefit: because coupling strengths are not adjusted globally, overflow or underflow issues related to the lifetime of c-bits become less of a concern, and they are a one-time problem to solve (for instance, past definitions of the c-bit have $dx_{i}/dt$ that grow exponentially as $\beta$ increases, thus overflows may occur at cold temperatures [38]).
We believe that the hybrid chaotic-probabilistic APT algorithm we propose here is widely applicable to a whole host of other phase-based oscillator Ising machines. However, the simplicity of the c-bit and its state-based (rather than phase) representation through explicit latches may still be more appealing.
VI.   Adaptive Parallel Tempering for Semiprime Factorization
Fig. 5: (a) p-bits are utilized to construct AND gates and full adders, which are the fundamental components for a multiplier. The product bits are clamped to a semiprime number. The system’s ground state finds the factors via invertible boolean logic [2]. (b) Energy profile of the 5 coldest replicas during hybrid chaotic-probabilistic APT. $E-E_{0}$ represents distance from ground state energy. (c) Time-to-solution vs. product size using a 10-bit $\times$ 10-bit factorizer for semiprime numbers. p-bit SA yields a slope of 0.441, while p-bit APT yields a slope of 0.438. c-bit SA exhibits a slope of 0.658, and hybrid APT has a slope of 0.420. (d) Average replica swap probabilities from factoring a specific semiprime number over 100 trials.
As a final example, we consider a 10-bit $\times$ 10-bit invertible multiplier circuit composed of p-AND gates and p-Full Adders [2, 5] (FIG. 5a). This multiplier is represented by an undirected network of $n=310$ p-bits with 1200 weighted connections, normalized such that $J_{ij}\in[-1,1]$. Of the $310$ spins in the system, 20 specific spins represent the product bits, while another set of 20 specific spins represents each of the two 10-bit factors. Using bipolar spin states ($m_{i}\in{\{-1,+1\}}$), traditional binary 00s are instead represented as negative spins $m_{i}=-1$. Clamping the multiplier’s 20-bit output to a semiprime number configures the system such that its ground state, $E_{0}$, solves for the correct prime factors. To accommodate semiprime numbers that are smaller than 20 bits, we pad the most significant bits with zeros. This scheme is referred to as invertible Boolean logic, in which p-bit based logic gates are run in reverse, similar in spirit to those used in memcomputing. [2, 59].
We employ identical parameters for both probabilistic APT and hybrid chaotic-probabilistic APT. Our APT pre-processing method finds a suitable $\beta$ profile (FIG. 5c). We consider $N=17$ system replicas, each running at a different inverse temperature $\beta$ for a duration of $t_{a}=10^{5}$ time steps. Probabilistic swap attempts are made between adjacent replica pairs once every 10 time steps, according to the Metropolis criterion (Equation 15). These parameters yield $Nt_{a}=17\times 10^{5}$ time steps of total computation summed across replicas.
During an APT simulation of $t_{a}=10^{5}$ time steps, if the coldest replica (highest $\beta$) detects the correct factors at any time step, the simulation is considered a successful solve (FIG. 5b). Only the coldest replica is considered to remove the effect of random search statistics which may play a role in small problem sizes. For our comparative study, optimizing the algorithms’ performance is not necessary.
We adopt a commonly used time-to-solution (TTS) metric [60]:
$\langle\text{TTS}\rangle=(Nt_{a})\left\langle{\frac{\ln{(1-0.99)}}{\ln{[1-p(Nt_{a})]}}}\right\rangle$ (17)
where $Nt_{a}$ is the total computational time summed across all replicas. The time-to-solution represents the amount of computation necessary to factor semiprimes with a 99% confidence level (see Appendix .3). For a given product size (number of bits), we consider 180 unique semiprime numbers. We attempt to factor each semiprime number 80 times. $p(Nt_{a})$ is measured as the number of successful solves divided by the total number of attempted factorizations.
We devise a similar scheme for simulated annealing for the sake of comparison. We anneal 17 system replicas in an embarrassingly parallel manner. Each replica runs for $t_{a}=10^{5}$ time steps, and $\beta$ is linearly increased from 0 to 10. In order to avoid random-find effects, we only check for the correct factors at the last 1000 time steps. If any of the 17 replicas finds the correct factors, the trial is considered a successful solve. For a given problem size (# bits), we again consider 180 unique semiprimes. We attempt to factor each semiprime 120 times. $p(Nt_{a})$ and time-to-solution are calculated in a fashion analogous to APT.
Once again, in order to rule out implementation-dependent differences in unit time between c-bits and p-bits, we perform a scaling analysis, where TTS vs. problem size is assumed to have exponential dependence: TTS $=A\exp{(kx)}$, where $x$ is the product size in bits, while $A$ and $k$ are constants.
Using this method of comparison, FIG. 5c shows that purely probabilistic APT exhibits remarkably similar algorithmic scaling to the newly introduced chaotic-probabilistic hybrid scheme. Moreover, they both exhibit better scaling than SA-based approaches. Pure chaotic SA, which has been commonly implemented [38], shows the worst scaling. Furthermore, both p-bits and c-bits observe a boundary effect as the problem size approaches 20 bits, where the time-to-solution actually decreases. This is a feature particular to the fixed 20-bit factorizer circuit: since numbers larger than 20 bits cannot be represented by the circuit, the solution space is reduced as we approach 20 bits. We exclude this data because it is uninformative for algorithmic scaling.
The two computing schemes not only share similar algorithmic scaling but also exhibit similarities in their underlying mechanisms. We consider a specific 11-bit semiprime number and conduct APT for $t_{a}=10^{5}$ time steps. We average over 100 random seeds. FIG. 5d shows the probability of the Metropolis criterion dictating a swap between a given pair of replicas. Remarkably, the hybrid APT scheme follows a similar qualitative distribution as the probabilistic APT scheme.
The choice of integer factorization as a benchmark problem is simply due to the challenging combinatorial nature it poses to stochastic local search algorithms [61]. Specialized factoring algorithms starting with trial division and more complex field sieve approaches exhibit superior algorithmic scaling. It is important to note, however, that the factoring representation we use in this work is based on the more general NP-hard Circuit SAT problem, for which no generic algorithmic speed up is known. As such, searching for scaling and prefactor improvements by dedicated hardware is highly desired.
VII.   Possible Hardware Realizations
Before we conclude, we make a few remarks regarding hardware implementations of chaotic networks. Chaotic bits have typically been implemented in digital and analog CMOS [45, 44]. Similar to CMOS implementations of p-bits, the main appeal of c-bit networks is in their asynchronous and massively parallel operation, provided that the synaptic input (Eq. 2) and the latching mechanism can be computed quickly to avoid detrimental parallel updates. We suspect that the initial randomization of oscillator phases and round-off errors in floating point software simulation are largely responsible for adding a layer of stochasticity to these networks. In hardware implementation, the needed noise can come from thermal noise in the environment. Beyond digital implementation, c-bits can make use of “natural” latches. For example, the physics of spin-torque switching of a non-volatile nanomagnet [62] can function as the latch: the magnet will switch only if a current threshold is reached, and the magnet will exhibit hysteresis until the current threshold changes sign, much like the billiard-ball metaphor used in c-bits.
A purported benefit of c-bits is the absence of explicit random numbers in their update rules. However, this claim requires qualifications. First, low-cost pseudo random number generators, such as LFSRs, do not significantly increase the complexity of implementing a p-bit. For example, an $n$-bit LFSR requires about $32\times n$ transistors to implement [58]. In contrast, constructing a c-bit digitally may be significantly more complex than this cost due to the necessity of an oscillator with a tunable duty cycle. Additionally, there may be “hidden” sources of randomness, such as floating point round-off errors in simulation or thermal noise in physical implementation.
Augmenting c-bits with explicit randomness in powerful tempering algorithms such as the one considered in this paper could offer further benefits. We believe that while the lack of random numbers may be an important c-bit advantage, comparisons must be made in concrete implementations in order to be meaningful (for example, against the most energy-efficient and compact hardware solutions of p-bits, using magnetic nanodevices [3]). Nevertheless, we believe that the similar performance of c-bits and p-bits across the wide variety of problems considered in this paper is very promising, providing valuable insights into the level of stochasticity required for effective probabilistic computing.
Conclusion
In this work, we evaluated and improved deterministic chaotic bit networks that combine billiard-like oscillators with latches. We demonstrated that augmenting these purely deterministic devices with stochasticity exhibits algorithmic scaling comparable to fully stochastic networks of p-bits. Additionally, we achieved better algorithmic scaling than common simulated annealing-based approaches for c-bits. Notably, the proposed algorithm is applicable to any other oscillator Ising machine.
Moreover, we have shown that c-bits approximately sample from the quantum Boltzmann law in a 1D TFIM model. Although c-bits exhibited qualitatively different critical scaling in the 2D Ising model and the 3D spin glass problem compared to the exact probabilistic bit formulations, their use in practical optimization problems remains promising. Combining stochastic p-bits and deterministic c-bits in asynchronous and massively parallel hardware implementations could create powerful domain-specific computers for combinatorial optimization and probabilistic sampling.
Acknowledgements
We gratefully acknowledge Sanaaya Lakdawala and Navid Anjum Aadit for useful initial discussions. We thank Nikhil Shukla for insights regarding Oscillator Ising Machines and Masoud Mohseni for useful discussions on the Kibble-Zurek mechanism. This work is partially supported by an Office of Naval Research Young Investigator Program grant, and a National Science Foundation CCF 2106260 grant. Use was made of computational facilities purchased with funds from the National Science Foundation (CNS-1725797) and administered by the Center for Scientific Computing (CSC). The CSC is supported by the California NanoSystems Institute and the Materials Research Science and Engineering Center (MRSEC; NSF DMR 2308708) at UC Santa Barbara.
Methods
Numerical simulation uses double precision (64-bit) in C++ programming language. Random numbers are generated using the Mersenne Twister engine. Spins (and initial billard states, $x_{i}$, for c-bits) are randomly initialized at the beginning of every SA and APT simulation. For c-bits, we employ Euler’s Method with a step of $\Delta t=0.1$ unless otherwise stated. At a given step, if the billiard $x_{i}$ is greater than or equal to $+1$ or less than or equal to $-1$, then $m_{i}$ is latched to that value.
For the 2D ferromagnetic Ising model, we obtain the critical energy experimentally. Using p-bits, we conduct $10^{6}$ Monte Carlo sweeps at constant critical temperature $T_{c}={2}/{\ln(1+\sqrt{2})}$. We record the average energy over the last $10^{5}$ sweeps. We take an ensemble average over 40 random seeds.
For the 3D spin glass instances, using p-bits, we conduct $10^{6}$ Monte Carlo sweeps at constant critical temperature $T_{c}=1.1$, a value found from a previous study [53]. For each of 300 problem instances, we record the average energy over the last $10^{5}$ sweeps.
.1 Appendix 1
In this Appendix, we describe the details of the Kibble-Zurek Mechanism (KZM) results shown in FIG. 3. Although we closely follow the arguments presented in Ref. [49], our purpose is to provide precise details about our own measurement procedure.
According to the scaling hypothesis in phase transitions, macroscopic quantities such as relaxation times and correlation lengths diverge at the critical point. These quantities typically exhibit a power-law relationship with respect to a tuning parameter, $\lambda$. In our context, $\lambda$ is the temperature, $T$, or its inverse, $\beta$. The power-law dependence of key quantities with respect to $\lambda$ around the critical point can usually be justified by notions of scale-invariance and self-similarity:
$\xi(\lambda)=\frac{\xi_{0}}{|\lambda_{c}-\lambda|^{\nu}}$ (18)
where $\xi$ is the correlation length, $\nu$ is the critical exponent for correlation lengths, and
$\tau(\lambda)=\frac{\tau_{0}}{|\lambda_{c}-\lambda|^{z\nu}}$ (19)
where $\tau$ is the relaxation time and $z\nu$ is the critical exponent for the relaxation time. Consider now a general annealing schedule, $\lambda(t)$ starting from an initial, $\lambda(t=0)=\lambda_{i}$ and ending at the critical $\lambda(t=t_{a})=\lambda_{c}$. Since the key quantities of interest diverge near the critical point, the system remains roughly in equilibrium until it approaches the critical point. The non-equilibrium dynamics of interest occur close to the critical point, thus any annealing schedule $\lambda(t)$ can be linearized around $\lambda_{c}$, which is reached at $t=t_{a}$:
$\lambda(t)=\lambda_{c}+\alpha(t-t_{a})+O[(t-t_{a})^{2}]$ (20)
where $\alpha$ is a key quantity representing the velocity of the anneal as the system approaches the critical point:
$\alpha=\left.\frac{d\lambda(t)}{dt}\right|_{t=t_{a}}$ (21)
As the system is annealed towards the critical point, the equilibration time increases. If the initial $\lambda$ is far away from the critical point, the system quickly relaxes to its equilibrium and can follow the annealing schedule in quasi-equilibrium or adiabatically. There comes a point, $\hat{t}$, when the remaining annealing time, $t_{a}-\hat{t}$ is precisely equal to the relaxation time of the system, $\tau[\lambda(\hat{t})]$. At this point, any further change in $\lambda$ that increases the relaxation time results in the system not being able to equilibrate by the end of the annealing. The idea of KZM, then, is to find this “freeze-out” boundary $\hat{t}$ and to relate non-equilibrium correlation lenghts to equilibrium correlations lengths via $\xi[\lambda(\hat{t})]$, which is then related to a density of topological defects via dimensional analysis. Based on this sketch, we proceed to find $\hat{t}$ by equating $|t_{a}-\hat{t}|$ to $\tau(\hat{\lambda})$ using Eq. 19 and Eq. 20. We obtain:
$\hat{t^{\prime}}=\bigg{(}\tau_{0}\alpha^{-z\nu}\bigg{)}^{\frac{1}{1+z\nu}}$ (22)
where $\hat{t^{\prime}}=t_{A}-\hat{t}$. We then find $\lambda(\hat{t^{\prime}})$ as:
$\lambda(\hat{t^{\prime}})=\lambda_{c}(1-\tau^{\prime}_{0}\alpha^{\frac{1}{1+z\nu}})$ (23)
where $\tau^{\prime}_{0}$ is a modified $\tau_{0}$ with unimportant constants that do not affect the exponents. We then obtain the correlation lengths at freeze-out from:
$\xi[\lambda(\hat{t})]=\epsilon^{\prime}_{0}\,\alpha^{{}^{\frac{-\nu}{1+z\nu}}}$ (24)
where $\epsilon^{\prime}_{0}$ is a modified $\epsilon_{0}$ with unimportant constants. Finally, $\xi$ is related to the density of topological defects through dimensional analysis of defect dimension $d$ and system dimension $D$:
$\frac{\xi^{d}}{\xi^{D}}=\frac{1}{\epsilon^{\prime(D-d)}_{0}}\,\alpha^{{}^{\left(\frac{\nu}{1+z\nu}\right)(D-d)}}$ (25)
Eq. 25 predicts a power-law relationship for the number of “defects” as a function of quench “velocity”, $\alpha$ in terms of equilibrium critical exponents. Defects in the context of Ising models are then relatable to measurable quantities such as kinks [51] and to the residual energy with similar power-law scaling laws. In this paper, we do not attempt to reproduce critical exponents for c-bits or p-bits, but simply observe the power-law dependence of c-bits and p-bits with respect to $\alpha$ in 2D ferro-Ising and 3D spin glass models.
.2 Appendix 2
Input: energy variance tolerance, number of chains, number of sweeps, number of samples, constant parameter $\alpha$
Output: problem-specific $\beta$ schedule for parallel tempering
$i\leftarrow 1$;
$\beta_{1}\leftarrow 0.5$;
Initialize chains to random spins;
while energy variance $\sigma_{E}$ is greater than tolerance do
       for each chain in parallel do
             Simulate p-bits using Eq. 1 and 2 for the specified number of Monte Carlo sweeps at constant inverse temperature $\beta_{i}$;
             Calculate variance $\sigma_{E}$ of energies sampled at the end of the simulation;
             Save the spin state $\{m\}$ of each chain as the initial condition for the next iteration;
       end for
      Calculate the mean energy variance across chains, $\langle\sigma_{E}\rangle$;
      $\beta_{i+1}\leftarrow\beta_{i}+\frac{\alpha}{\langle\sigma_{E}\rangle}$;
       $i\leftarrow i+1$;
end while
Algorithm 2 Adaptive Parallel Tempering Preprocessing with p-bits
We use the following parameters for APT preprocessing: 100 parallel chains, $10^{4}$ Monte Carlo sweeps, 1000 samples (taken at the end of $10^{4}$ sweeps), and $\alpha=1.25$. Energy variance tolerance is equal to half of the smallest magnitude weight $J_{ij}$, and the initial value of $\beta$ is set to 0.5. We have used a similar preprocessing method in our previous work [57].
.3 Appendix 3
To derive our time-to-solution performance metric (Eq. 17), we start from:
$[1-p(Nt_{a})]^{k}=1-0.99$ (26)
where $p(Nt_{a})$ is the probability of finding the solution using $N$ system replicas each running for time $t_{a}$. The left-hand side of the equation represents the probability of failing $k$ times in a row, and the right-hand side is the complement of the desired success probability, 0.99. We solve for $k$ through algebraic manipulation. The time-to-solution is then:
$Nt_{a}\times k$ (27)
which yields our definition of time-to-solution (Eq. 17). This can be thought of as the total computational time of the algorithm, $Nt_{a}$, multiplied by the average number of attempts before success.